Keller Company makes two models of battery-operated boats, the Sandy Beach and the Rocky River. Basic production information follows:

Sandy Beach Rocky River
Direct materials cost per unit $ 19.30 $ 26.60
Direct labor cost per unit 14.80 19.10
Sales price per unit 83.40 105.00
Expected production per month 1,200 units 910 units

      

Keller has monthly overhead of $11,104, which is divided into the following cost pools:

Setup costs $ 2,400
Quality control 5,766
Maintenance 2,938
Total $ 11,104

The company has also compiled the following information about the chosen cost drivers:

Sandy Beach Rocky River Total
Number of setups 11 29 40
Number of inspections 180 285 465
Number of machine hours 1,300 1,300 2,600

Required:
1.
Suppose Keller uses a traditional costing system with machine hours as the cost driver. Determine the amount of overhead assigned to each product line. (Do not round intermediate calculations and round your final answers to the nearest whole dollar amount.)

Overhead assinged

Sandy Beach

Rocky River

Total Overhead

Please explain how you got this and not just answer. Thank you and will give you a thumbs up
Answer #1

Overhead rate under traditional costing = Estimated overheads/Estimated machine hours

= 11,104/2600

= $4.27076923 per hour

Overhead assigned = Driver consumed*Overhead rate

Sandy Beach = Machine hours*Rate per hour

= 1300*4.27076923 = $5,552

Rocky River = 1300*4.27076923 = $5,552

Total overhead = $11,104